Factorise the following expressions
 10a^2 -15b^2 +20c^2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

10a^2-15b^2+20c^2
10a^2= 2×5×a×a
- 15b^2= -1×3×5×b×b
20c^2= 2×2×5×c×c
Common factor of 10a^2, 15b^2and 20c^2is 5
10a^2-15b^2+20c^2= 5(2a^2-3b^2+4c^2)

Thanks